OpenStudy (anonymous):

square root of y^15? Simplify

OpenStudy (anonymous):

(y^15)^(1/2)=y^(15x(1/2))=y^(15/2)

OpenStudy (anonymous):

not sure what you mean

OpenStudy (anonymous):

\[y^{15/2} \]cannot be simplified.

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

actually it was y^7 outside of the square root of y, thank you anyways

OpenStudy (anonymous):

From Mathematica:\[\text{Simplify}\left[y^7y^{1/2}\right]\to y^{15/2} \]Takes a little less ink to print I guess.